OpenStudy (anonymous):

Find the constant of variation for the relationship f(x)= 40x. A. f(x) B. x C. 20 D. 40

OpenStudy (anonymous):

f(x) = 40x means that the constant of variation is 40

OpenStudy (anonymous):

ok. thanks What about this one?

OpenStudy (anonymous):

Find the constant of variation for the relationship shown in the following table: 01_09_LQ_2.gif A. 2 B. 3 C. 4 D. 1

OpenStudy (anonymous):

table says x|1|2|3|4 y|3|6|9|12|

OpenStudy (anonymous):

Okay here is how you find this hun. Look at the top line, "X" it is going "1, 2, 3, 4" So how much is it going by to get the next number?

OpenStudy (anonymous):

@noneyabusiness ? you there?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

+1?

OpenStudy (anonymous):

That's right! So x has a constant of +1 now what about Y ?

OpenStudy (anonymous):

+3

OpenStudy (anonymous):

Yup! So that means that for every x (1), you add 3 to y. So if the next number on the line x was 5, y would be 15. See how it works?

OpenStudy (anonymous):

Is the constant 3?

OpenStudy (anonymous):

Yes. Y has a constant of +3

OpenStudy (anonymous):

so answer B

OpenStudy (anonymous):

yay! you are really smart!

OpenStudy (anonymous):

Thank you :D

OpenStudy (anonymous):

you are welcome

OpenStudy (anonymous):

Any more you nee help with?

OpenStudy (anonymous):

yes please

OpenStudy (anonymous):

If f(x) varies directly with x, and f(x) = 10 when x = 4, write the direct linear variation equation. A. f(x)=10x B. f(x=4x C. f(x)=5/2x D. f(x)=2/5x Is it b?

OpenStudy (anonymous):

Okay, let me figure it out, I will post the answer in just a sec

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Okay, it is C, f(x)= 5/2x

OpenStudy (anonymous):

i see how you did it too

OpenStudy (anonymous):

Okay cool, You need help with any more?

OpenStudy (anonymous):

nope! thanks though

OpenStudy (anonymous):

You are absolutely welcome!